# Interaction of water surface waves with periodic and quasiperiodic cylinder arrays

### v1

@@ Abstract
Inspired by transformation optics and photonic crystals, this paper presents a computational investigation into the interaction between water surface waves and array waveguides of cylinders with multiple previously unexplored lattice geometries, including, for the first time, quasiperiodic geometries. Extending beyond conventional square and hexagonal periodic arrays, transformation optics has opened up entirely new opportunities to investigate water wave propagation through arrays based on quasiperiodic lattices, and quasiperiodically arranged vacancy defects. Using the linear potential flow open-source code Capytaine, missing element and -scaled Fibonacci square lattices, the Penrose lattice, hexagonal lattice and Amman-Beenker lattice are investigated. The existence of band gaps for all arrays is observed. An hexagonal lattice with vacancy defects transmits the least energy. Bragg diffraction consistent with rotational symmetry is observed from all arrays. Waves will distort significantly to achieve resonance with arrays, supporting transformation-based waveguides. The possible uses include adaptation to more versatile waveguides with applications such as offshore renewable energy and coastal defence.
I Introduction
@@
Control of water wave propagation is an essential capability for multiple applications in coastal and offshore engineering. The power in surface waves can pound cliffs and beaches – contributing to the increasing global problem of coastal erosion – or it can be harnessed to provide energy to help break society’s reliance on fossil fuels. Improving our ability to manage and harness this energy is a problem of paramount importance, and has been the focus of much work over recent years [41, 39, 40, 36, 34, 15, 16, 17, 33, 38, 23]. Indeed, free-surface water waves (disturbances of water) have been comprehensively studied analytically [24], numerically [19], and through observations both in controlled physical experiments and field measurements [6, 7]. It has been shown that water surface waves share many of the same features as waves in other areas of physics (including fundamental processes such as reflection, refraction, diffraction, etc.). However, water waves also exhibit unique behaviour including shoaling (change of wave height) and breaking (collapse of waves) when they propagate into progressively shallower water. While the increasing emphasis on extracting energy from ocean waves and coastal protection has expanded our knowledge and understanding, new research is still needed, particularly to improve our understanding of complex wave interactions and the exploitation of wave energy devices.
One avenue of exploration over the last few decades has been the investigation of the interactions between water surface waves and arrays of cylinders, particularly in the context of offshore engineering. This has received renewed interest with the recent focus on offshore renewable energy and alternative coastal protection schemes. Interestingly, but not surprisingly, it has been shown that the problem of large arrays is analogous to electron propagation through crystals, leading to the application of the Bloch theorem to infinite and finite arrays of cylinders, and consequently the derivation of the band structure of these arrays [25]. Similarly, Ha et al. generated transmission spectra of finite square and hexagonal arrays for wave propagation along high-symmetry lattice directions, and found complete band gaps in each case [14].
@@
In addition to periodic lattices built on crystalline symmetries, quasiperiodic lattices may also be constructed that have no translational symmetry but higher-order rotational symmetries, such as the pentagonal Penrose tiling [28] and the octagonal Ammann-Beenker tiling [3]. Quasicrystals (quasiperiodic crystals) based on these lattices are the appropriate analogue of periodic crystals [31, 21]. The higher orders of rotational symmetry available in quasiperiodic structures are attractive because they offer the same behaviours that periodic structures do, such as band gaps, but with greater isotropy and hence less angular selectivity [30].
The similarities observed in electron and water surface wave propagation have also motivated analogies to other wave physics, such as transformation optics. Transformation optics is a relatively new field devoted to tailoring the propagation of electromagnetic waves which has had remarkable breakthroughs for superlenses and invisibility devices [27]. Particular success has been found in constructing photonic crystals: metamaterials with structures on length scales similar to the wavelength of the radiation. Photonic crystals and photonic quasicrystals have both been shown to exert deep influence on electromagnetic wave propagation [18, 13]. In this case, modifying the permittivity and the permeability of the propagation medium manipulates the propagation of the waves. A similar field is transformation acoustics, in which the Young’s modulus of the propagation medium is modified, with similarly useful results [8]. The analogous properties which affect propagation of water waves are gravity and depth. Quasicrystalline metamaterials have been shown to have fractal transmission spectra with several large gaps [11]. Despite the clear analogies between transformation optics and water wave propagation, such lattice arrangements of cylinder arrays and scatterers are yet to be fully investigated for water waves.
@@
With respect to water surface waves, and in the long-wavelength regime (, where is the wavelength and is a characteristic crystal dimension), it has been shown that the presence of an array of vertical cylinders alters the refractive index via modification of the effective gravity within the array, allowing construction of lenses and prisms [15]. This approach has also been shown experimentally to result in superlensing [17]. Moreover, analogous to the manipulation of permittivity/permeability in photonic materials, manipulating the depth via modification of either the floor or the free surface can lead to the full range of transformation optics effects [37, 29]. For example, the refractive index can be varied continuously by applying a gradient to the floor [35], or a band gap can develop via interaction with a periodically drilled floor [16]. However, there has been limited investigation into the exploitation of rotational symmetries within the lattice array structures.
@@
Here, inspired by the similarities with transformation optics, we explore the interaction between monochromatic water surface waves and small () arrays of vertical surface-piercing cylinders. These cylinders are arranged in translational and rotational symmetries previously uninvestigated and, in particular, we measure their capacity to allow or stop wave propagation in the near-Bragg-resonance regime (). Our focus is on the detailed structure of the arrays, which include periodic, quasiperiodic, and defected periodic geometries with the resulting effects on water surface wave propagation.
@@
The paper is structured as follows. First, we introduce the arrays that are investigated, and characterise them in terms of their symmetries. Then, we describe the wave modelling and numerical setup using the open-source software Capytaine [2, 4], before presenting the results from the simulations. We finish with a discussion on future work and conclusions.
II Lattices
@@
Our interest lies in exploring the wave propagation properties of arrays with varying rotational symmetries and periodic/quasiperiodic structures. The designs investigated are shown in Figure 1, encompassing archetypes and simple modifications thereof of 4–, 6–, 8–, and 10–fold rotational symmetries, where they are generally referred to by their -fold symmetry as f. The finer details of their structure and nomenclature are discussed below. In an effort to study the effects of geometry independently of other characteristics of the array, where possible, the most common nearest neighbour distance is fixed at m. Cylinders are placed at the lattice points (section III.2) to calculate the filling fraction (FF) defined as the area occupied by cylinders divided by the total area of the array (see Figure 2(b)). According to the work of Hu et al.[15], the refractive index is then . Number of elements, FF and are given for each array in Table 1. For the purposes of comparison both periodic and aperiodic arrays are investigated.
@@
II.1 Periodic arrays
The two simplest classes of periodic lattice are 4–fold (square) and 6–fold (hexagonal), arrays based on which are shown in Figure 1 and labelled as 4f and 6f respectively, where the subscript s denotes ‘simple’. All continuous lattices in two-dimensional space must have 2–fold rotational symmetry.
II.2 Quasiperiodic arrays
Missing-element square Fibonacci lattice
Array 4f is based on a square lattice with elements missing in an quasiperiodic fashion. The missing elements are chosen based on the binary Fibonacci sequence (LSLLSLS..), where the individual elements of the sequence are either 2a (S, short) or 3a (L, long).
-scaled square Fibonacci lattice
Array 4fτ is a square lattice of elements placed according to 2 perpendicular Fibonacci sequences, where is equal to the golden ratio . As there are times as many segments as segments in each (infinite) sequence, the in this case is , and the FF is accordingly high [22].
Hexagonal lattices based on the lattice
Array 6f is based on a periodic hexagonal lattice with elements removed to produce a quasiperiodic arrangement, similar to 4f. The lattice points decorate the vertices of the lattice identified by Coates et al. [10], where the short and long edges of the rhombic and hexagonal tiles are both set to 1 m [9]. Array 6f is also produced using the lattice with tile edges set to 1 m, but instead uses the centre points of the tiles.
Penrose lattice
Array 10f is formed from the vertices of a Penrose P3 lattice [28], which is perhaps the most well-known quasiperiodic geometry. The lattice is composed of two types of rhombi, with internal angles (thin) and (fat), arranged according to matching rules such as to give the tiling an overall 10–fold symmetry, although it is often described as pentagonal.. The rhombus edge length is chosen to be 1 m. This again leads to elements being separated by less than 1 m, here, along the short diagonal of the thin rhombi.
III Methods
III.1 Numerical domain
The array is placed at the centre of a 40 m 40 m square grid, with a grid resolution of 512 512 and a resultant cell size of 7.8 7.8 cm. The waves are incident from the left hand side. They impinge on the array, are scattered, and we observe the steady state response of the water wave properties in the domain.
The water is set to infinite depth so that the waves can be considered deep water waves. The arrays themselves are composed of fixed 0.2 m radius, 10 m long, 13-gon prism approximations to cylinders, with 5 m submerged. As noted, the most common separation present in the array is set to 1 m. Wave heights are scaled from unity in Capytaine: any wave height or motion amplitude can be retrieved by multiplying the result by the desired value [2], and the results are valid in the linear regime.
III.2 Water wave modelling with Capytaine
@@
IV Results and discussion
Here, we present and discuss the results of our simulations in terms of relevant wave phenomena: diffraction, blocking, and resonance, before directly comparing the periodic and quasiperiodic arrays in terms of their performance.
IV.1 Diffraction
In general, all of the arrays produce strong diffraction when the Bragg condition is satisfied by one or more separations in the array. The diffracted beams are directional and diffuse, which is consistent with the small number of scatterers in the arrays.
We use the Bragg law to extract the apparent interplanar distance of the arrays when the wavefront is normal to the array. Close to the Bragg angles for a given lattice, constructive interference may occur for a range of frequencies, with the diffracted beam sweeping a concomitant range of angles. Therefore, it is appropriate to use the Bragg law to extract only when a lattice has one or more easily identified directions of symmetry. Figure 4 shows this analysis for the 4f, 6f, 8f, and 10f arrays, where a black arrow indicates the incoming and diffracted wave direction and angle associated with the rotational symmetry of the array (). For the simple square array, is found to be close to the expected value of , and for the simple hexagonal array, it is close to twice the expected value of , as our frequency range is too low to observe primary Bragg reflection. The deviation from the expected values is consistent with the finite size of the array and the finite radius of the scatterers. For example, for array 4f, scattering is evident from the entire range of lattices between those defined by the innermost points and outermost points of the cylinders in the array. In this array, the minimum number of planes at 45∘ is 11. This gives an approximate % error from the limits of , given by (.
The 4f, 6f, and 6f arrays show the same broad behaviour as the simple structures with this analysis. The 8f and 10f arrays have well-defined scattering angles, which permits the extraction of an apparent ‘interplanar’ distance, shown in the figure as .
The -scaled 4–fold array 4fτ, not shown here, has ambiguous behaviour: although it has 4–fold symmetry, it never strongly diffracts from a set of planes at 45∘ as expected. However, weak diffraction at multiple wavelengths yields values for of 1.00 m and 1.39 m.
IV.2 Blocking

### v2

@@ Abstract
Inspired by transformation optics and photonic crystals, this paper presents a computational investigation into the interaction between water surface waves and array waveguides of cylinders with multiple previously unexplored lattice geometries, including, for the first time, quasiperiodic geometries. Extending beyond conventional square and hexagonal periodic arrays, transformation optics has opened up entirely new opportunities to investigate water wave propagation through arrays based on quasiperiodic lattices, and quasiperiodically arranged vacancy defects. Using the linear potential flow open-source code Capytaine, missing element and -scaled Fibonacci square lattices, the Penrose lattice, hexagonal lattice and Amman-Beenker lattice are investigated. The existence of band gaps for all arrays is observed. A hexagonal lattice with vacancy defects transmits the least energy. Bragg diffraction consistent with azimuthal rotational symmetry is observed from all arrays. Bragg resonance causes reflection from arrays, resulting in multiple Bloch band gaps. Away from Bragg resonance, waves will distort significantly to achieve periodic relationships with arrays, supporting transformation-based waveguides. The possible uses include adaptation to more versatile waveguides with applications such as offshore renewable energy and coastal defence.
keywords:
waveguides , Capytaine , quasiperiodic , transformation opticsPACS:
47.11.-j ,PACS:
61.44.Br ,MSC:
52C23 ,MSC:
76B15[inst1]organization=Jeremiah Horrocks Institute for Mathematics, Physics and Astronomy, University of Central Lancashire,city=Preston, postcode=PR1 2HE, country=United Kingdom
[inst2]organization=Department of Physics, University of Liverpool,city=Liverpool, postcode=L69 3BX, country=United Kingdom
[inst3]organization=School of Engineering and Computing, University of Central Lancashire,city=Preston, postcode=PR1 2HE, country=United Kingdom
[inst4]organization=School of Engineering, University of Manchester,city=Manchester, postcode=M13 9PL, country=United Kingdom
Inspired by transformation optics, quasiperiodic array water waveguides investigated
Propagation through arrays supports a transformation approach to waveguide design
Near 100% redirection of wave energy achieved with minimal arrangements
Bragg resonance and Bloch band gaps in quasiperiodic lattices identified
Bragg diffraction consistent with rotational symmetry
1 Introduction
The vast potential of the revolution in optical waveguide engineering known as ‘transformation optics’ has stimulated optics research for the past two decades [1, 2]. Combined with contemporary developments in manufacturing capability, particular success has been found in constructing photonic crystals: metamaterials with structures on length scales similar to the wavelength of the radiation. By modifying the permittivity and the permeability of the propagation medium, photonic crystals have been shown to exert deep influence on electromagnetic wave propagation [3, 4], leading to the demonstration of superlenses and invisibility devices [1]. A similar field is transformation acoustics, in which the Young’s modulus of the propagation medium is modified, with similarly useful results [5].
Inspired by transformation optics and photonic crystals, in this article we present an investigation into the interaction of water surface waves with ‘metamaterials’ consisting of arrays of vertical cylinders in crystalline and other arrangements.
@@ 1 Introduction
Improving our ability to manage and harness the energy in ocean waves is a problem of paramount importance, and has been the focus of much work over recent years [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Free-surface water waves (disturbances of water) have been comprehensively studied analytically [17], numerically [18], and through observations both in controlled physical experiments and field measurements [19, 20]. It has been shown that water surface waves share many of the same features as waves in other areas of physics (including fundamental processes such as reflection, refraction, diffraction, etc.). However, water waves also exhibit unique behaviour including shoaling (change of wave height) and breaking (collapse of waves) when they propagate into progressively shallower water. While the increasing emphasis on extracting energy from ocean waves and coastal protection has expanded our knowledge and understanding, new research is still needed, particularly to improve our understanding of complex wave interactions and the exploitation of wave energy devices.
There are several approaches to investigating the interaction of water waves with structures. In the predominant approach, which started before simulation via computation became mainstream, efforts are focused on identifying those systems which lend themselves to analytical solutions. The majority of these systems are 2-dimensional problems, being the propagation dimension and being the depth. The obstacles are modelled as variations in the propagation medium as functions of and , including floating horizontal, vertical or otherwise mathematically defined bars, plates and 1-dimensional periodic structures [21, 22, 23], fixed versions of similar structures [24], or similar structures applied to the sea bed [25, 26].
Parallel behaviour in the interactions between water waves and periodic structures and between x-rays and crystals was noticed early on – in particular, the advantageous Bragg resonance, in which waves are reflected from a periodic array and a Bloch frequency band gap is formed in which waves cannot propagate [27, 28]. In arrangements which are symmetrical in the propagation direction, zero-reflection wave modes, in which waves are able to propagate freely through the structure, are also possible [25, 29]. These phenomena in crystals underpin x-ray and electron diffraction, and the understanding of the behaviour of electrons in solids, including the band gap in semiconductors and conduction in metals [30].
Where analytical solutions were possible, they were developed, including of 3-dimensional systems of infinite 2-dimensional arrays of modulations of the propagation medium such as floating discs, based on square (4-fold) and triangular/hexagonal (6-fold) lattices [31]. Arrays of surface-piercing cylinders were also investigated, with periodic boundary conditions (PBC) for infinite arrays [32, 11, 16] and without PBC for finite arrays [33, 34, 35]. To the best of our knowledge, all infinite and finite periodic arrangements exhibit Bloch band gaps. The relationship between a 2-dimensional array and the propagation direction, however, is no longer implicit; waves with different propagation directions experience different lattice periodicities and different transmission properties, including band gap. The variability in periodicity is a function of the -fold rotational symmetry of the lattice on which the array is based.
@@
In addition to periodic lattices based on crystalline symmetries, quasiperiodic lattices may also be constructed that have no translational symmetry but higher-order rotational symmetries (higher values of ), such as the pentagonal Penrose tiling (10-fold) [36] and the octagonal Ammann-Beenker tiling (8-fold) [37]. Quasicrystals (quasiperiodic crystals) based on these lattices are the appropriate analogue of periodic crystals [38, 39]. The higher orders of rotational symmetry available in quasiperiodic structures are attractive because they offer the same behaviours that periodic structures do, such as band gaps, but with greater isotropy and hence less angular selectivity [40], supporting, in principle, manipulation of waves travelling in different directions. Quasicrystalline metamaterials have been shown to have fractal transmission spectra with several large gaps [41].
Regarding the comparison with transformation optics (in which the relative permittivity and permeability are modified), the analogous properties which affect propagation of water waves are gravity and depth. Despite the clear analogies between transformation optics and water wave propagation, such lattice arrangements of cylinder arrays and scatterers are yet to be fully investigated for water waves.
@@
With respect to water surface waves, and in the long-wavelength regime (, where is the wavelength and is a characteristic crystal dimension, defined as the lattice parameter for periodic arrays), it has been shown that the presence of an array of vertical cylinders alters the refractive index via modification of the effective gravity within the array, allowing construction of lenses and prisms [11]. This approach has also been shown experimentally to result in superlensing [13]. Moreover, analogous to the manipulation of permittivity/permeability in photonic materials, manipulating the depth via modification of either the floor or the free surface can lead to the full range of transformation optics effects [42, 43, 44].
Generally, modifying the density of a lattice of elements of given permeability/permittivity [Young’s modulus] causes optical [acoustic] waves to propagate in the direction of the density modifications. The situation for water surface waves is essentially the same.
Regarding analytical approaches, as mentioned, the application of PBC allows the derivation of transmission functions for linear waves through arrays [45]. If the problem is restricted to two dimensions, as is possible for infinite depth or floor-mounted cylinders, it is possible to derive transmission functions even for finite arrays [16]. However, as suggested by the name, PBC cannot describe quasiperiodic geometries of any kind. In principle PBC may be applied to higher-dimensional space, from which a quasiperiodic two- or three-dimensional solution may be obtained (using the cut-and-project method [46, 47]). Such a solution preserves the general behaviour expected from an infinite quasiperiodic lattice, but interpreting the implications in real space is non-trivial and a source of unwanted complexity.
Regarding computational approaches, any conceivable system may be investigated using fully-non-linear potential flow (FNLPF) calculations, but these calculations are extremely expensive in terms of time and computing resources. Alternatively, the popular boundary-element-method (BEM) may be used to investigate arbitrary three-dimensional problems within the fully linear regime (including those that can be simplified with PBC, with concomitant increases in efficiency).
Here, we demonstrate the use of the Capytaine [48] BEM python package to investigate a three-dimensional problem: finite arrays of finite-length cylinders in infinite-depth water. This problem, despite its apparent simplicity, is intractable via any other method except FNLPF calculations, and is thus ideal to demonstrate the validity of the BEM method to study arrays with arbitrary properties.
The robustness of the approach with respect to arbitrary arrays is of great importance for the development of the field of water wave array waveguides. The versatility of the approach enables the study of an arbitrary number of optimization parameters and thus, in principle, the tailoring of a waveguide to a particular situation.
@@
We explore the interaction between monochromatic water surface waves and small () arrays of vertical surface-piercing cylinders. These cylinders are arranged in translational and rotational symmetries previously uninvestigated and, in particular, we measure their capacity to allow or stop wave propagation in the near-Bragg regime (). Our focus is on the detailed structure of the arrays, which include periodic, quasiperiodic, and defected periodic geometries with the resulting effects on water surface wave propagation.
@@
The paper is structured as follows: first, we introduce the arrays that are investigated, and characterise them in terms of their symmetries. Then, we describe the wave modelling and numerical setup using the open-source software Capytaine [48, 49], before presenting the results from the simulations. We finish with a discussion on future work and conclusions.
2 Lattices
@@
Our interest lies in exploring the wave propagation properties of arrays with varying rotational symmetries and periodic/quasiperiodic structures. The designs investigated are shown in Figure 1, encompassing archetypes and simple modifications thereof of 4–, 6–, 8–, and 10–fold rotational symmetries, where they are generally referred to by their -fold symmetry as f. The finer details of their structure and nomenclature are discussed below. In an effort to study the effects of geometry independently of other characteristics of the array, where possible, the most common nearest neighbour distance is fixed at m. A cylinder is placed at each lattice point (section 3.2). The filling fraction (FF) is defined as the area occupied by cylinders divided by the total area of the array (see Figure 2(b)). Each infinite lattice has a calculable FF. However, in keeping with the desire to deal only with properties of finite arrays, here FF is calculated directly by summing the cylinder cross-sectional areas and dividing by the area of the bounding box of the array. According to the work of Hu et al. [11], the refractive index is then . Number of elements, FF and are given for each array in Table 1. For the purposes of comparison both periodic and aperiodic arrays are investigated. We note that the derivation of Hu et al. contains no reference to the array geometry at any stage, implying its validity for both periodic and aperiodic arrays. We also note that it is derived in the long wavelength limit; however, the correspondence principle implies that long-wavelength-regime phenomena will also apply in the Bragg regime we are operating in, although they will be dominated by other effects.
2.1 Periodic arrays
The two simplest classes of periodic lattice are 4–fold (square) and 6–fold (hexagonal), arrays based on which are shown in Figure 1 and labelled as 4f and 6f respectively, where the subscript s denotes ‘simple’. All continuous lattices in two-dimensional space must have even-number , due to their indistinguishability under rotational inversion [50].
2.2 Quasiperiodic arrays
Missing-element square Fibonacci lattice
Array 4f is based on a square lattice with elements missing in an quasiperiodic fashion. The missing elements are chosen based on the binary Fibonacci sequence (LSLLSLS..), where the individual elements of the sequence are either a (S, short) or 2a (L, long). The sequence is indicated in the figure.
-scaled square Fibonacci lattice
Array 4fτ is a square lattice of elements placed according to 2 perpendicular Fibonacci sequences, one of which is indicated in the figure, where is equal to the golden ratio . As there are times as many segments as segments in each (infinite) sequence, the in this case is , and the FF is accordingly high [51].
Hexagonal lattices based on the lattice
Array 6f is based on a periodic hexagonal lattice with elements removed to produce a quasiperiodic arrangement, similar to 4f. The lattice points decorate the vertices of the lattice identified by Coates et al. [52], where the short and long edges of the rhombic and hexagonal tiles are both set to 1 m [53]. Array 6f is also produced using the lattice with tile edges set to 1 m, but instead uses the centre points of the tiles. In these cases, the relationship to the Fibonacci sequence is not trivial, so it is not indicated. Further details can be found in refs. [52, 53].
Ammann-Beenker lattice
Penrose lattice
Array 10f is formed from the vertices of a Penrose P3 lattice [36], which is perhaps the most well-known quasiperiodic geometry. The lattice is composed of two types of rhombi, with internal angles rad (thin) and rad (fat), arranged according to matching rules. Though it is often described as pentagonal, its indistinguishability under rotational inversion gives this lattice overall 10–fold symmetry [50]. The rhombus edge length is chosen to be 1 m. This again leads to elements being separated by less than 1 m, here, along the short diagonal of the thin rhombi, with the separation being equal to m.
@@ 2.2 Quasiperiodic arrays
3 Methods
3.1 Numerical domain
The array is placed at the centre of a 40 m 40 m square grid, with a grid resolution of 512 512 and a resultant cell size of 7.8 7.8 cm. The calculations result in linear solutions to a linear problem. This means that the grid resolution of the water surface does not affect the calculation or the results, in contrast with the effects of the cell size in FNLPF simulations. The linear solution is sampled at each point in the grid, and the grid size set as the lowest power of 2 that preserves the smallest details observed in the data. Convergence testing of the grid resolution for 8 values in the range 64 to 1024 yields a constant (L2 error norm vs ) gradient of 0.175, where is equal to the computational domain edge length (40 m) divided by the grid resolution. For well-converged results, the gradient is not greater than 1. The gradient in our test is much smaller than 1, so the results are converged. The gradient in this case is entirely due to the improvement expected from an increase in sampling resolution, and cannot therefore decrease further.
The waves are incident from the left hand side. They impinge on the array, are scattered, and we observe the steady state response of the water waves in the domain. The water is set to infinite depth so that the waves can be considered deep water waves. The arrays themselves are composed of fixed 0.2 m radius, 10 m long, 13-gon prism approximations to cylinders, with 5 m submerged. The number of sides is chosen to maximise resolution whilst keeping within memory constraints, and also to avoid any rotational symmetries in common with the arrays. As noted, the most common separation present in the array is set to 1 m. Wave heights are scaled from unity in Capytaine: any wave height or motion amplitude can be retrieved by multiplying the result by the desired value [48], and the results are valid in the linear regime.
3.2 Water wave modelling with Capytaine
@@
4 Results and discussion
There are several facets to the interaction between the array waveguides and the impinging water surface waves. We begin with a description of the ‘blocking’ behaviour of the waveguides; that is, the capacity of a waveguide to prevent wave energy from reaching the leeward side of the array. This behaviour depends, to varying degrees according to the various waveguides, on the primary Bragg resonance.
We then describe other Bragg diffraction behaviour. As we are dealing with three-dimensional simulations of 2-dimensional arrays, we have the opportunity to observe non-primary Bragg resonances characteristic of the rotational symmetry of the waveguides. These cause effective redirection of the wave energy.
We finish with a description of the transmission of wave energy through the waveguides, which is accompanied and enabled by the establishment of a periodic relationship between wave and waveguide.
4.1 Blocking